OpenStudy (anonymous):

Solve the equation. (9x/7)-x=(x-28)/21

OpenStudy (anonymous):

x = -28/5

OpenStudy (anonymous):

i need workkkk ..

jhonyy9 (jhonyy9):

(9x/7)-x=(x-28)/21 (9x-7x)/7 =(x-28)/21 3(2x)=x-28 6x=x-28 5x= -28 x= -28/5